OpenStudy (anonymous):

square root of x minus 7 end root plus 5 equals 11

OpenStudy (psymon):

You want to isolate the square root portion of your problem, which can be done by subtracting 5 from both sides, leaving us with: \[\sqrt{x-7}= 6\] Now we want to raise both sides to the 2nd power. This allows us to eliminate the square root we have. \[(\sqrt{x-7})^{2}= 6^{2}\] After the root is gone you just isolate x. Think you can go from there?

OpenStudy (anonymous):

Yes, thank you :)